# HardFlow: Hard-Constrained Sampling for Flow-Matching Models via Trajectory Optimization

### v1

@@ I Introduction
Despite some empirical success, projections during sampling suffer from several fundamental limitations. First, these methods are intrinsically conservative as they enforce pathwise feasibility, a condition that is unnecessarily restrictive since only the terminal state (i.e., the sample of interest) must satisfy the constraint. The intermediate iterates are more of algorithmic artifacts. By imposing constraints on these transient states, projection during sampling prematurely prunes the search space, which can prevent the generative process from discovering higher-quality solutions that would otherwise be accessible. Although per-step projection can be interpreted as projected gradient ascent on the data log-likelihood [15], this view relies on Langevin-style sampling with specific noise and step-size schedules and does not generally extend to arbitrary SDE discretizations or the ODE-based samplers used in flow matching. Second, these projection-based methods solely address constraint satisfaction and are not designed to optimize a reward function, whereas an ideal solution should achieve both feasibility and optimality.
@@ IV Problem Formulation
Concretely, given $\bar{x}_{0}\sim p_{0}$, we seek a per-sample refinement procedure that steers each trajectory during ODE integration to a terminal state $\tilde{x}_{1}$ such that (i) the hard constraints $h(\tilde{x}_{1})\leq 0$ are satisfied, (ii) the cost $C(\tilde{x}_{1})$ is typically low, and (iii) collectively, over draws of $\bar{x}_{0}$, the resulting terminal distribution $x_{1}\sim\tilde{\mu}$ remains close to the nominal distribution $\bar{\mu}$.
@@ V Algorithm Design
Remark 2.
Although we perform a series of transformations on Problem 2 to obtain Problem 6, there is no relaxation of terminal feasibility. For any $x$, we have $\mathcal{M}_{t_{N}}^{\theta}(x)=x$, $(\mathcal{M}_{t_{N}}^{\theta})^{-1}(x)=x$, and $\mathcal{F}_{t_{N}}^{\theta}(x)=x$. Therefore $h(x_{N})\leq 0$ is enforced exactly through the last subproblem in all receding-horizon formulations.

### v2

@@ I Introduction
Despite some empirical success, projections during sampling suffer from several fundamental limitations. First, these methods are intrinsically conservative as they enforce pathwise feasibility, a condition that is unnecessarily restrictive since only the terminal state (i.e., the sample of interest) must satisfy the constraint. The intermediate iterates are more of algorithmic artifacts. By imposing constraints on these transient states, projection during sampling prematurely prunes the search space, which can prevent the generative process from discovering higher-quality solutions that would otherwise be accessible. Although per-step projection can be interpreted as projected gradient ascent on the data log-likelihood [15], this view relies on Langevin-style sampling with specific noise and step-size schedules and does not generally extend to arbitrary SDE discretizations or the ODE-based samplers used in flow matching. Second, these projection-based methods solely address constraint satisfaction and are not designed to optimize a reward function. In practical applications, however, it is often desirable not only to enforce hard constraints but also to optimize rewards to improve sample quality. An ideal method should jointly address both aspects and produce feasible samples of high quality.
@@ IV Problem Formulation
Concretely, given $\bar{x}_{0}\sim p_{0}$, we seek a per-sample refinement procedure that steers each trajectory during ODE integration to a terminal state $\tilde{x}_{1}$ such that (i) the hard constraints $h(\tilde{x}_{1})\leq 0$ are satisfied, (ii) the cost $C(\tilde{x}_{1})$ is typically low, and (iii) collectively, over draws of $\bar{x}_{0}$, the resulting terminal distribution $\tilde{x}_{1}\sim\tilde{\mu}$ remains close to the nominal distribution $\bar{\mu}$.
@@ 6  Solve the optimization problem
It is worth emphasizing that, although we perform a series of transformations on Problem 2 to obtain Problem 6, there is no relaxation of terminal feasibility. More precisely, we have the following result.
Proposition 1.
Suppose the feasible set $\mathcal{S}=\{x\mid h(x)\leq 0\}$ is nonempty. The output sample $x_{N}$ of Algorithm 1 satisfies the hard constraints $h(x_{N})\leq 0$.
Proof.
At the final iteration i=N1i=N-1, let $\widehat{x}_{N}^{*}$ denote a solution to (11). Then $h(\widehat{x}_{N}^{*})\leq 0$. Also note that $t_{N}=1$. The boundary conditions of the affine scheduler $(\alpha_{t},\beta_{t})$ at $t=1$ are $\alpha_{1}=1$ and $\beta_{1}=0$.
Therefore, for i=N1i=N-1, we can simplify the update rule in Line 1 of Algorithm 1 as
$x_{N}=\alpha_{1}\widehat{x}_{N}^{*}+\beta_{1}\frac{-\dot{\alpha}_{1}\bar{x}_{N}+\alpha_{1}v_{1}^{\theta}(\bar{x}_{N})}{\alpha_{1}\dot{\beta}_{1}-\dot{\alpha}_{1}\beta_{1}}=\widehat{x}_{N}^{*}.$ (12)
Consequently, $h(x_{N})=h(\widehat{x}_{N}^{*})\leq 0$, which shows that the output sample $x_{N}$ satisfies the hard constraints. ∎
@@
Remark 3.
The constrained optimization problem (11) can be solved using any suitable method, depending on the specific downstream application. In general, it may be solved numerically with constrained optimization algorithms, such as sequential quadratic programming or interior-point methods. Moreover, if the cost $C(\cdot)$ and constraints $h(\cdot)$ possess particular structures (e.g., a quadratic cost and linear constraints), specialized solvers or even closed-form solutions may be available.